OpenStudy (anonymous):

If f(x) = integral (3,x) sqr root (1+t^3) dt, find (f^1)'(0)

OpenStudy (anonymous):

\[f(x) = \int\limits_{3}^{x} \sqrt{1+t ^{3}} dt\] find \[(f ^{-1})'(0)\]

OpenStudy (anonymous):

\[f(x) = \int\limits_{3}^{x} \sqrt{1+x ^{3}} dx\] \[f(x) = (1+x^3)^{1/2}\] \[f'(x) = \frac{ 1 }{ 2 }(1+x ^{3})^{^{-1/2}} (3x ^{2})\]

OpenStudy (slaaibak):

f'(x) = sqrt(1+x^3) so f'(0= 1

OpenStudy (anonymous):

is that setting up: putting the 0 from (f^-1)' \[0=(1+x^3)^{\frac{ 1 }{ 2 }}\]

OpenStudy (slaaibak):

no. so f'(0)= 1 meaning sqrt(1+0^3) = 1

OpenStudy (anonymous):

you just want to sent the equation equal to 0?

OpenStudy (anonymous):

oops set

OpenStudy (slaaibak):

bro, you asked me to find f'(0) and I found f'(0), what are you on about?

OpenStudy (slaaibak):

you know evaluating a function at zero and setting it equal to zero are two different things?

OpenStudy (anonymous):

correct

OpenStudy (slaaibak):

so you gave f(x). I found f'(x) and then I evaluated it at zero (f'(0)) That's the question, isn't it?

OpenStudy (anonymous):

Yes we evaluated f'(x) at zero and we ant to put that into \[(f ^{-1})'(0) = \frac{ 1 }{ f'(f ^{-1}(0) }\] Correct? Or am I too tired and my brain is done for the day.

OpenStudy (anonymous):

\[= \frac{ 1 }{ f'(1) }\]

OpenStudy (slaaibak):

oh! I completely overlooked the inverse over there, sorry

OpenStudy (anonymous):

yeah those lovely inverses

OpenStudy (anonymous):

do we need to back track some for the steps done for doing f'(0) or are we good?

OpenStudy (slaaibak):

so yeah, you solve 0 = f(x) then x=3 do you agree? Then f^-1(0) = 3 ? or alternatively, f(3) = 0

OpenStudy (anonymous):

ok

OpenStudy (slaaibak):

then f'(3) = sqrt(1+3^3) = sqrt(28)

OpenStudy (anonymous):

dang I didn't need to figure that ugly f'(x)?

OpenStudy (slaaibak):

you did. Well, I did. \[f'(x) ={d \over dx} \int\limits_{3}^{x} \sqrt{1+t^3}dt = \sqrt{1+x^3}\]

OpenStudy (slaaibak):

it's kinda obvious if you think about it. you should get the same answer back, since you are integrating and then differentiating again. BUT there is a chain rule adjustment that has to be made if there are more complex bounds

OpenStudy (anonymous):

I wish these derivative inverses made more sense to me. I am pretty confident in my inverse solving. But these inverse derivatives are killing me.

OpenStudy (slaaibak):

haha, just remember the methods/formulas, do it step by step and you'll be fine

OpenStudy (anonymous):

Thanks for your help slaaibak. Have a great night.

OpenStudy (slaaibak):

No problem. Thanx, you too.

OpenStudy (anonymous):

ok can some one work this problem for me from 1st step to last step. I see you plug x in for t. and 3 in for x but what about the (f^-1)(0) what does that fit in thanks